OpenStudy (anonymous):

Express 5/21 as a decimal...please help fan and medal!!

OpenStudy (anonymous):

0.23809523809

OpenStudy (anonymous):

whoa haha thanks you!!

OpenStudy (anonymous):

wait umm

OpenStudy (anonymous):

theres lines on top of the answers. kinda weird

OpenStudy (whpalmer4):

yes, that means the string of digits under the line repeats...

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

okay hole on

OpenStudy (anonymous):

0.238095

OpenStudy (anonymous):

oh okay thank you!!

OpenStudy (whpalmer4):

You can see that here, as I've showed you to 100 places: 0.23809523809523809523809523809523809523809523809523809523809523809523\ 80952380952380952380952380952381

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

can you help me on one more last one?

OpenStudy (whpalmer4):

Some other common fractions that do that: Anything divided by 7 Anything divided by 3 ("anything" excludes numbers not divisible by that number, of course!) \[\begin{array}{cc} \frac{1}{7} & 0.142857142857142857142857142857 \\ \frac{2}{7} & 0.285714285714285714285714285714 \\ \frac{3}{7} & 0.428571428571428571428571428571 \\ \frac{4}{7} & 0.571428571428571428571428571429 \\ \frac{5}{7} & 0.714285714285714285714285714286 \\ \frac{6}{7} & 0.857142857142857142857142857143 \\ \end{array}\]

OpenStudy (whpalmer4):

11s\[\begin{array}{cc} \frac{1}{11} & 0.0909090909090909090909090909091 \\ \frac{2}{11} & 0.181818181818181818181818181818 \\ \frac{3}{11} & 0.272727272727272727272727272727 \\ \frac{4}{11} & 0.363636363636363636363636363636 \\ \frac{5}{11} & 0.454545454545454545454545454545 \\ \frac{6}{11} & 0.545454545454545454545454545455 \\ \frac{7}{11} & 0.636363636363636363636363636364 \\ \frac{8}{11} & 0.727272727272727272727272727273 \\ \frac{9}{11} & 0.818181818181818181818181818182 \\ \frac{10}{11} & 0.909090909090909090909090909091 \\ \end{array}\]

OpenStudy (whpalmer4):

(if the digit pattern changes on the very last digit, that's due to rounding, not the pattern changing)

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

can you help me on a last one?

OpenStudy (whpalmer4):

well, only if you post it :-)

OpenStudy (anonymous):

ohh ok hole on

OpenStudy (anonymous):

hold

OpenStudy (anonymous):

Write 4.48 as the quotient of two integers.

OpenStudy (anonymous):

336/75

OpenStudy (anonymous):

i dont have that for one of my answer choises...

OpenStudy (whpalmer4):

\[4.48 = 4 + \frac{48}{100} \]right?

OpenStudy (whpalmer4):

\[4 = \frac{400}{100}\] \[\frac{400}{100} + \frac{48}{100} =

OpenStudy (anonymous):

try simplyfing

OpenStudy (whpalmer4):

ackthphth. \[\frac{400}{100} + \frac{48}{100}=\]

OpenStudy (anonymous):

??????????????

OpenStudy (anonymous):

what?

OpenStudy (whpalmer4):

\[4.48 = \frac{448}{100}\]does it not?

OpenStudy (anonymous):

no i dont have that for my answr choises

OpenStudy (whpalmer4):

Sigh. Well, whether it is an answer choice or not, 4.48 is the quotient of the integers 448 and 100. You can try reducing the fraction until you find a fraction that is on your list of answer choices. Or you can try dividing all of the answer choices and see which one is equal to 4.48.

OpenStudy (anonymous):

okay thanks though

OpenStudy (whpalmer4):

\[\frac{448}{100} = \frac{224}{50} = \frac{112}{25}\]Any of those appear on your list?

OpenStudy (anonymous):

its okay nvm